Find the value of a for which the polynomial x^5-ax²-ax+1has −1 as a root with multiplicity at least 2.
1
Expert's answer
2020-03-05T16:06:02-0500

ANSWER: a=-5

EXPLANATION.

"P_5(x):=x^5-ax^2-ax+1=(x^5+1)-ax(x+1)" "\quad { { P }_{ 4 }(x)=\frac { { P }_{ 5 }(x) }{ x+1 } ={ x }^{ 4 }{ -x }^{ 3\quad }+{ x }^{ 2 } }-x+1-ax" . "P_4(-1)=0" , since "x=-1" is the root of the second multiplicity of the polynomial "P_5(x)" . At the same time, "P_4(-1)=a+5" .Therefore "a+5=0" or "a=-5" .

VERIFICATION: If "a=-5" , then "P_5(x)=x^5+5x^2+5x+1" , "\quad \frac { { P }_{ 5 }(x) }{ { \left( x+1 \right) }^{ 2 } } ={ x }^{ 3 }-2{ x }^{ 2 }+3x+1"